# OmniRetarget: Interaction-Preserving Data Generation for Humanoid Whole-Body Loco-Manipulation and Scene Interaction

### v1

@@ I Introduction
    Successful zero-shot sim-to-real transfer of proprioceptive RL policies on a physical humanoid, demonstrating a diverse set of scene-interaction tasks, including a long, agile sequence of object carrying, platform climbing, jumping, and rolling.
@@ VI-A Different Sources of Human Motion Data
Human motion datasets contain rich pose and shape information but differ in format as well as in the body shapes and heights of demonstrators. Thus, we need to convert these inputs into a consistent representation, typically a time series of global 3D keypoint positions $\{p^{\text{source}}_{0:T,i}\}$, and scale the motions to account for differences between human demonstrators and the target robot, before solving the kinematics.
@@
Different retargeting pipelines use different strategies to handle these formats. We detail these preprocessing steps below, denoting the human demonstrator’s pose as $q^{\text{demo}}_{t}$, the SMPL forward model for the $i$-th keypoint as $M_{i}$, the original demonstrator shape as $\beta^{\text{source}}$, and the demonstrator’s $i$-th keypoint position as $p^{\text{demo}}_{t,i}$.

### v2

@@ I Introduction
    Successful zero-shot sim-to-real transfer of proprioceptive RL policies on a physical humanoid, demonstrating a diverse set of scene-interaction tasks, including a long, agile sequence of object carrying, platform climbing, jumping, rolling and wall-flipping.
@@ V-A Whole-Body Scene-Object-Interaction
We additionally showcase a high-dynamic wall-flip motion [^1] The motion is acquired from https://actorcore.reallusion.com/3d-motion?asset=parkour-tic-tac-backflip . An IMU capable of measuring angular rates above $15$ rad/s is required for this motion. in Fig. 5. The robot completes the full flip in approximately $0.5$ second, reaching a peak angular velocity of $15$ rad/s. Unlike the human foot, which can flex at the arch to maintain contact and generate friction, the robot foot is rigid. As a result, it must align more closely to the wall to achieve sufficient contact area and friction. To account for this physical difference and give RL more freedom to learn this skill, we relaxed the termination condition during RL training by increasing the end-effector position error threshold to $0.5$ meter (compared to $0.25$ meter used in other motions) and removed the foot joint orientation tracking term from the reward function. All other components of the tracking objective remain consistent with other motions. The trained policy is robust and achieves a $5/5$ success rate in our real-world experiments.
Figure 5: Hardware results showing a high-dynamic wall-flip motion. The robot reaches a maximum linear velocity of $3.5$ m/s and a peak angular velocity of $15$ rad/s.
@@ VI-A Different Sources of Human Motion Data
Human motion datasets contain rich pose and shape information, but they differ both in data format and in the physical attributes (e.g., height, body proportions) of the demonstrators. To make them compatible across different sources and suitable for retargeting, we need to convert these inputs into a consistent representation, typically a time series of global 3D keypoint positions $\{p^{\text{source}}_{0:T,i}\}$. This process must account for differences between human demonstrators and the target robot.
@@
Different retargeting pipelines use different strategies to handle these formats. We detail these preprocessing steps below, denoting the human demonstrator’s pose as $q^{\text{demo}}_{t}$, the SMPL forward model for the $i$-th keypoint as $M_{i}$, the original demonstrator shape as $\beta^{\text{demo}}$, and the demonstrator’s $i$-th keypoint position as $p^{\text{demo}}_{t,i}$.